### v1

@@ Abstract
Accurately predicting blood glucose (BG) levels of ICU patients is critical, as both hypoglycemia (BG < 70 mg/dL) and hyperglycemia (BG > 180 mg/dL) are associated with increased morbidity and mortality. We develop the Multi-source Irregular Time-Series Transformer (MITST), a novel machine learning-based model to forecast the next BG level, classifying it into hypoglycemia, hyperglycemia, or euglycemia (70-180 mg/dL). The irregularity and complexity of Electronic Health Record (EHR) data, spanning multiple heterogeneous clinical sources like lab results, medications, and vital signs, pose significant challenges for prediction tasks. MITST addresses these using hierarchical Transformer architectures, which include a feature-level, a timestamp-level, and a source-level Transformer. This design captures fine-grained temporal dynamics and allows learning-based data integration instead of traditional predefined aggregation. In a large-scale evaluation using the eICU database ((200,859 ICU stays across 208 hospitals), MITST achieves an average improvement of 1.7% ($p<0.001$) in AUROC and 1.8% ($p<0.001$) in AUPRC over a state-of-the-art baseline. For hypoglycemia, MITST achieves an AUROC of 0.915 and an AUPRC of 0.247, both significantly higher than the baseline’s AUROC of 0.862 and AUPRC of 0.208 (p-value < 0.001). The flexible architecture of MITST allows seamless integration of new data sources without retraining the entire model, enhancing its adaptability in clinical decision support. Although this study focuses on predicting BG levels, MITST can easily be extended to other critical event prediction tasks in ICU settings, offering a robust solution for analyzing complex, multi-source, irregular time-series data.
@@ 1 Introduction
Conventional methods of glucose management often rely on periodic measurements and reactive treatments, which may not be optimal for maintaining glycemic control in a dynamic and high-risk setting like the ICU. Accurate prediction of glucose levels can facilitate timely interventions and improve patient outcomes. Traditionally, this problem has been approached using regular time-series forecasting techniques (e.g., ARMA, ARIMA) [19, 20, 21] or various learning algorithms (e.g., linear models, random forests) [22, 23, 24]. However, a significant challenge is their limitations in managing the highly irregular nature of electronic health record (EHR) data, as these methods typically require regular time interval-based measurements for accurate forecasting. Also, contemporary learning algorithms applied to this task often rely on aggregating the irregular EHR data to fit the model’s needs [25]. Unfortunately, predefined aggregation schemes tend to overlook crucial temporal information, which can contain significant insights into a patient’s condition. For instance, aggregating data over fixed intervals can obscure critical patterns and trends that occur over shorter or varying timescales.
@@ 4.2 Joint representation
The input $\mathbf{z}(m)$ is first normalized with a LayerNorm operation and then passed through a linear transformation $\text{Linear}_{1}(d^{\prime}_{m},\,d^{\prime}_{m}\,\times\,mult\,\times\,2)$, where $mult$ controls the expansion factor of the hidden layer. A GEGLU [32] activation function is applied to introduce non-linearity. The result is then projected to the target dimensionality $d_{j}$ via a second linear transformation $\text{Linear}_{2}(d^{\prime}_{m}\,\times\,mult,\,d_{j})$. The final output $\mathbf{u}(m)$ represents the source $m$ in the shared space. This process is applied independently to each source, yielding the set of shared representations $\{\mathbf{u}(m)\}_{m=1}^{M}$. This step ensures that the transformed data from each source can be effectively combined and compared in the subsequent steps.
@@ 7 Conclusions and future work
While the current study demonstrates the effectiveness of supervised learning in predicting BG levels, future work could explore the application of self-supervised learning techniques. This approach would be particularly useful in scenarios where labeled data is scarce or difficult to obtain. Additionally, future work could explore the use of multi-task learning [38] to further enhance MITST’s versatility. Multi-task learning would enable the model to simultaneously learn multiple related tasks—such as predicting other vital signs or clinical outcomes—by sharing representations across tasks, improving generalization and efficiency. Also, from an ethical standpoint, it is essential to recognize potential biases within the model due to the underlying dataset, which may not fully represent all ICU patient populations. To mitigate this risk, future studies should prioritize fairness and explainability [39], ensuring that the model does not disproportionately benefit or harm any specific subgroup of patients.

### v2

@@ Abstract
Accurately predicting blood glucose (BG) levels of ICU patients is critical, as both hypoglycemia (BG < 70 mg/dL) and hyperglycemia (BG > 180 mg/dL) are associated with increased morbidity and mortality. This study presents a proof-of-concept machine learning framework, the Multi-source Irregular Time-Series Transformer (MITST), designed to predict blood glucose (BG) levels in ICU patients. Unlike existing approaches that rely on manual feature engineering or are limited to a small number of Electronic Health Record (EHR) data sources, MITST demonstrates the feasibility of integrating diverse clinical data (e.g., lab results, medications, vital signs) and handling irregular time-series data without predefined aggregation. MITST employs a hierarchical architecture of Transformers, comprising feature-level, timestamp-level, and source-level components, to capture fine-grained temporal dynamics and enable learning-based data integration. This eliminates the need for traditional aggregation and manual feature engineering. In a large-scale evaluation using the eICU database (200,859 ICU stays across 208 hospitals), MITST achieves an average improvement of 1.7% ($p<0.001$) in AUROC and 1.8% ($p<0.001$) in AUPRC over a state-of-the-art baseline. For hypoglycemia, MITST achieves an AUROC of 0.915 and an AUPRC of 0.247, both significantly outperforming the baseline. The flexible architecture of MITST allows seamless integration of new data sources without retraining the entire model, enhancing its adaptability for clinical decision support. While this study focuses on predicting BG levels, MITST can easily be extended to other critical event prediction tasks in ICU settings, offering a robust solution for analyzing complex, multi-source, irregular time-series data.
@@ 1 Introduction
Conventional methods of glucose management often rely on periodic measurements and reactive treatments, which may not be optimal for maintaining glycemic control in a dynamic and high-risk setting like the ICU. A significant challenge of current literature on BG prediction task is their limitations in managing the highly irregular nature of electronic health record (EHR) data, as these methods typically require regular time interval-based measurements for accurate forecasting. Also, contemporary learning algorithms applied to this task often rely on aggregating the irregular EHR data to fit the model’s needs [19]. Unfortunately, predefined aggregation schemes tend to overlook crucial temporal information, which can contain significant insights into a patient’s condition. For instance, aggregating data over fixed intervals can obscure critical patterns and trends that occur over shorter or varying timescales.
@@
The rest of the paper is organized as follows. In Section 2, we review the related work relevant to our research. Section 3 defines our problem, followed by Section 4, which presents the data material used in this study. In Section 5, we describe our proposed method in detail. In Section 6, we provide our experimental setup and results. We discuss our findings in Section 7, while concluding and outlining future directions in Section 8.
2 Related work
BG prediction methods can be broadly classified into two categories: physiology-based and data-driven approaches [21, 22, 23]. Physiology-based methods model BG dynamics using compartments such as meal absorption, insulin activity, exercise, and individual glucose metabolism [21]. However, our focus is on data-driven approaches. For a detailed review of physiology-based methods, we refer interested readers to [23]. Data-driven methods for BG prediction fall under the broader category of clinical time-series forecasting. These datasets are often irregularly sampled, making traditional time-series models unsuitable without pre-processing steps like data imputation or window-based aggregation. Imputation-based techniques [24] aim to convert irregular time-series into regular ones for use with standard forecasting models. For example, Lipton et al. [25] applied forward- and back-filling methods on 13 clinical time-series features and used long short-term memory (LSTM) networks to predict diagnoses. Similarly, Harutyunyan et al. [26] combined imputation with indicator variables and evaluated an LSTM on four clinical tasks. Alternatively, window-based approaches aggregate task-specific features from time-series segments. For example, Harutyunyan et al. [26] used windowed features with logistic regression for ICU mortality prediction.
Some studies have attempted to directly model ITS data by learning a continuous-time model from discrete and irregular observations [27, 28]. For example, [28] models ITS using multiple Gaussian processes (GPs) in the lower level of a hierarchical system and captures the transitions between GPs by using a linear dynamical system. With advancements in deep learning, ITS-focused methods have emerged. Baytas et al. [29] proposed Time-Aware LSTM (T-LSTM), which adjusts hidden states based on time gaps. Bai et al. [30] used an attention mechanism to handle irregular time-intervals between clinical events for diagnosis prediction. Tipirneni et al. [31] introduced a Transformer-based framework for ITS, learning triplet embeddings (feature name, value, timestamp) to predict mortality.
Unlike most existing methods, MITST inherently handles ITS data without requiring time-series-level imputation. Instead, imputation is applied only at the feature level to fill missing values within timestamp-specific vectors, preserving the raw irregularity of the time-series data. Two notable exceptions are T-LSTM [29] and the Transformer-based model by Tipirneni et al. [31], which also adapt to the irregularity of time intervals directly within their architectures. T-LSTM models patient data as a sequence of discrete (i.e., non-consecutive) visits, requiring larger datasets across multiple patient encounters, whereas MITST focuses on data from a single ICU stay, making it applicable to scenarios with fewer data points. Also, while Tipirneni et al. evaluated their Transformer-based model on mortality prediction—a simpler task compared to next BG level forecasting—MITST is tested on a more complex task and eliminates the need for manual feature selection, which was required in their approach. Moreover, while most prior works use a limited number of EHR sources (e.g., diagnosis or procedure codes), MITST integrates a broader range of clinical data (e.g., vital signs, lab results, medications). This allows the model itself to determine the the most relevant combination of data sources for a specific clinical task, eliminating the need for predefined feature engineering that is common in traditional ITS methods.
It is worth noting that Transformers, though dominant in natural language processing, are less mature in time-series forecasting [32]. Traditional machine learning models, such as random forests, XGBoost, or ARIMA combined with engineered features, remain competitive in this domain. This also holds for large language models (LLMs) in medical applications [33]. Thus, we compare MITST to the state-of-the-art random forest model developed by Zale et al. [19], which employs carefully designed features for next BG prediction.
@@ 5.2 Joint representation
The input $\mathbf{z}(m)$ is first normalized with a LayerNorm operation and then passed through a linear transformation $\text{Linear}_{1}(d^{\prime}_{m},\,d^{\prime}_{m}\,\times\,mult\,\times\,2)$, where $mult$ controls the expansion factor of the hidden layer. A GEGLU [39] activation function is applied to introduce non-linearity. Note that the same activation function is also used in all Transformer layers due to its better performance [39]. The result is then projected to the target dimensionality $d_{j}$ via a second linear transformation $\text{Linear}_{2}(d^{\prime}_{m}\,\times\,mult,\,d_{j})$. The final output $\mathbf{u}(m)$ represents the source $m$ in the shared space. This process is applied independently to each source, yielding the set of shared representations $\{\mathbf{u}(m)\}_{m=1}^{M}$. This step ensures that the transformed data from each source can be effectively combined and compared in the subsequent steps.
@@ 8 Conclusions and future work
While the current study demonstrates the effectiveness of supervised learning in predicting BG levels, future work could explore the application of self-supervised learning techniques. This approach would be particularly useful in scenarios where labeled data is scarce or difficult to obtain. Additionally, future work could explore the use of multi-task learning [45] to further enhance MITST’s versatility. Multi-task learning would enable the model to simultaneously learn multiple related tasks—such as predicting other vital signs or clinical outcomes—by sharing representations across tasks, improving generalization and efficiency. Another direction for future work is the application of large language models (LLMs) in EHR data analysis, utilizing their capability to extract insights from unstructured clinical notes and improve multi-source data integration. Also, from an ethical standpoint, it is essential to recognize potential biases within the model due to the underlying dataset, which may not fully represent all ICU patient populations. To mitigate this risk, future studies should prioritize fairness and explainability [46], ensuring that the model does not disproportionately benefit or harm any specific subgroup of patients.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgements
Mehdizavareh and Khan were supported by the Novo Nordisk Foundation [grant number NNF22OC0072415]. Cichosz was supported by i-SENS, Inc (Seoul, South Korea).